OpenStudy (rissyroo118):

hELP Solve the equation. |k + 9| = 4 {–13, –5} {–5, 5} {–5, 13} {all real numbers greater than or equal to –13 and less than or equal to –5}

rvc (rvc):

k+9=+4 k+9=-4

OpenStudy (rissyroo118):

sorry that didnt help

rvc (rvc):

solve for it

rvc (rvc):

a+b=d a=d-b

OpenStudy (rissyroo118):

answer d thank you actually hwlped alot

rvc (rvc):

wait k=-4-9

OpenStudy (rissyroo118):

What expression is equivalent to ? |a|≤4? a ≥ –4 and a ≤ 4 a ≤ –4 or a ≥ 4 a ≥ –4 and a ≤ 4 a ≤ –4 and a ≥ 4

rvc (rvc):

listen -4-9=-13 ok

rvc (rvc):

@whpalmer4

OpenStudy (rissyroo118):

itd be a -5 & -13

rvc (rvc):

Yes!

OpenStudy (rissyroo118):

thank you {: can you help with a few more?